OpenStudy (anonymous):

5x greater than or equal -10 Explain

OpenStudy (paxpolaris):

\[5x \ge-10\]

OpenStudy (paxpolaris):

divide both sides by 5

OpenStudy (anonymous):

Explain please... I want to have a understand this inequality better.

OpenStudy (anonymous):

This means that 5 times an unknown number x is greater than or equal to -10. Divide by 5 on both sides of the equation and you will get \[x \ge-2\]. This means the value of the unknown number is either 2 or something greater than 2. It can be 2, 3, 4, 5, or even 2.0001. The possibilities are infinite so you just express the answer as \[x \ge-2\]